# SPA: Efficient User-Preference Alignment against Uncertainty in Medical Image Segmentation

### v1

@@ Abstract
Medical image segmentation data inherently contain uncertainty, often stemming from both imperfect image quality and variability in labeling preferences on ambiguous pixels, which depend on annotators’ expertise and the clinical context of the annotations. For instance, a boundary pixel might be labeled as tumor in diagnosis to avoid under-assessment of severity, but as normal tissue in radiotherapy to prevent damage to sensitive structures. As segmentation preferences vary across downstream applications, it is often desirable for an image segmentation model to offer user-adaptable predictions rather than a fixed output. While prior uncertainty-aware and interactive methods offer adaptability, they are inefficient at test time: uncertainty-aware models require users to choose from numerous similar outputs, while interactive models demand significant user input through click or box prompts to refine segmentation. To address these challenges, we propose SPA, a segmentation framework that efficiently adapts to diverse test-time preferences with minimal human interaction. By presenting users a select few, distinct segmentation candidates that best capture uncertainties, it reduces clinician workload in reaching the preferred segmentation. To accommodate user preference, we introduce a probabilistic mechanism that leverages user feedback to adapt model’s segmentation preference. The proposed framework is evaluated on a diverse range of medical image segmentation tasks: color fundus images, CT, and MRI. It demonstrates 1) a significant reduction in clinician time and effort compared with existing interactive segmentation approaches, 2) strong adaptability based on human feedback, and 3) state-of-the-art image segmentation performance across diverse modalities and semantic labels.
@@ 1 Introduction
Deep learning-based medical image segmentation has achieved remarkable advancements over the past decade [28, 32, 6]. However, existing approaches often fall short when applied to real-world dynamic clinical environments. A critical challenge is to handle the inherent uncertainties in medical images [13]. A single medical image may have multiple different valid segmentations, depending on the labeling criteria under a specific medical context. For instance, in glioma detection from brain CT scans, it is often preferred to include surrounding tissue than risk missing part of the tumor [2]. While in radiation therapy for low-grade glioma, undersegmentation is used to protect sensitive brain tissue from excess radiation damage [27]. Therefore, developing adaptive methodologies that align segmentation uncertainties with specific labeling preferences for different clinical needs is essential.
@@
Figure 1: (a) Existing uncertainty-aware models suffer from choosing among numerous similar-looking candidates, making the process labor-intensive and time-consuming. (b) Interactive segmentation models lack the ability to incorporate image uncertainty and rely heavily on pixel-level interactions, which require a substantial amount of time and effort. (c) Our uncertainty-aware interactive segmentation model, SPA, efficiently achieves preference-aligned segmentation by incorporating medical image uncertainties and human interactions. Clinicians are presented with one recommended prediction and a few distinct segmentation candidates that capture uncertainty, allowing them to select the one best aligned with their clinical needs. If the user is unsatisfied with the recommended prediction, the model learns from the user selection, adapts itself, and presents users a new set of candidates. Our approach minimizes user interactions and eliminates the need for painstaking pixel-wise adjustments compared to conventional interactive segmentation models.
@@
Existing uncertainty-aware approaches [17, 29, 3, 22] represent segmentation uncertainty to users by generating numerous stochastic predictions (Fig. 1). However, these approaches require clinicians to painstakingly choose from countless similar-looking candidates, making the process overwhelming and time-consuming. Furthermore, since these models cannot incorporate human feedback for adjustments, there is a risk that none of the stochastic predictions may provide a satisfactory segmentation.
@@
Incorporating human feedbacks, often in the form of visual prompts, into the segmentation has achieved significant progress [31, 7, 24, 18]. However, most of such interactive segmentation approaches are unable to incorporate uncertainty. Additionally, existing visual prompts like clicks and bounding boxes are often cumbersome in real world, as they demand pixel-level interaction input from the users (Fig. 1).
@@
To address the above challenges, we believe the most effective solution is to use a small number of distinct candidates to represent segmentation uncertainty, allowing users to make a straightforward multi-choice selection. In this paper, we introduce SPA, a new paradigm for efficient Segmentation Preference Alignment with uncertainty in medical image segmentation (Fig. 1). SPA presents image uncertainties by generating multiple segmentations. Instead of overwhelming clinicians with numerous similar-looking predictions as in conventional uncertainty-aware segmentations, our model offers one recommended prediction and four representative correction proposals per iteration. Once the user selects an option, SPA rapidly adapts to the corresponding labeling preference, allowing convergence to the desired segmentation within only a few iterations. Experimental validation demonstrates that clinicians can segment 35% more images with 39% fewer iterations compared to previous interactive models, highlighting SPA’s potential for real-world clinical applications. In summary, our contributions are:
  • We introduce SPA, the first framework specifically designed to align user preferences with segmentation uncertainty, minimizing both interaction time and clinician effort.
  • We propose a preference distribution using mixture Gaussians to represent image uncertainties, enabling rapid adaptation to user preferences with minimal iterations.
  • We develop a multi-choice interaction mechanism that captures user preferences from modeled uncertainties, achieving preference-aligned predictions while reducing clinician interaction effort.
@@
  • We compare SPA with deterministic, uncertainty-aware, and interactive SOTA methods. SPA achieves superior results measured in Dice Score across multi-clinician annotated datasets such as REFUGE2, LIDC-IDRI and QUBIQ. It also consistently outperforms interactive methods within fewer rounds of user interactions.
@@
Figure 2: Overall framework of SPA. The inference process comprises of two steps: Preference-aware Segmentation and Preference Adaptation with Human Feedback. At timestamp $j$, SPA takes the input image $\mathbf{x}$, an interaction embedding $e_{p}^{(j)}$, and samples $\{\mathbf{z_{n}}^{(j)}\}_{n=1}^{N}$ drawn from the preference distribution $p^{(j)}_{\theta}(z)$ to generate $N$ segmentation predictions. These predictions are combined into a aggregated preference-aware prediction $\mathbf{\hat{y}_{app}}^{(j)}$ via majority vote. If the clinician does not approve $\mathbf{\hat{y}_{app}}^{(j)}$, SPA generates $K$ correction proposals $\{\mathbf{\Delta_{r_{k}}}^{(j)}\}_{k=1}^{K}$ in the adaptation step. The clinician selects the preferred proposal and the preference distribution is updated to $p^{(j+1)}_{\theta}(z)$ based on this choice. This process iterates until the segmentation meets clinical satisfaction.
@@ 2 Related Work
Medical Image Segmentation Medical image segmentation involves predicting a dense pixel-wise map to delineate structures such as lesions, tumors, and anatomical regions. It plays an indispensable role in disease diagnosis and treatment planning [39, 35, 26]. Deep learning has significantly enhanced segmentation accuracy and efficiency across various clinical applications [5, 34], with many models demonstrating remarkable success [28, 32, 6]. However, these models often fail to address the inherent uncertainty in medical images, as multiple plausible segmentations may exist for a single image [13]. To improve medical image segmentation, it is crucial to develop methods that address these uncertainties and identify the most suitable segmentation for the specific clinical context.
Uncertainty-aware Segmentation Uncertainty in medical images cannot be reduced by adding more data or using more complex models [19]. Techniques such as model ensembling, label sampling [15], and multi-head strategies [11] attempt to address this uncertainty by generating a range of potential predictions that reflect different user preferences [40, 14]. Probabilistic segmentation methods, including ProbUNet [20], PhiSeg [3], CM-Net [37], and MRNet [16] explicitly model the posterior distribution of parameters or predictions to capture uncertainty. However, these approaches often generate multiple predictions, requiring clinicians to review each one individually. Moreover, the predictions may not perfectly match the specific clinical context and some techniques rely on prior knowledge of the clinician’s expertise, complicating their adoption in clinical practice. In contrast, a more efficient solution is to incorporate human interactions, allowing clinicians to refine the segmentations through direct interaction to better align with the clinical context.
Interactive segmentation Interactive segmentation is an iterative process where automated segmentation results are refined through user input until they meet the desired output. Previous methods [33, 30, 18] have effectively modified predictions based on user interaction at the pixel level, with many achieving notable success in the medical domain [9, 25, 36, 38]. However, these interactive models often fail to address the inherent uncertainty in medical images. Additionally, they may require numerous iterations to align with the user preference for specific clinical context, given the existence of multiple valid segmentations. An interactive model capable of incorporating image uncertainties and learning from user interactions would reduce the number of iterations required to achieve the desired outcomes.
@@ 3 Methodology
3.1 Framework Overview
@@
Medical image segmentation faces challenges due to variations in labeling preferences, often influenced by clinical guidelines for downstream analysis. These variations can result in multiple plausible annotations for a single image, highlighting the need for adaptable segmentation models that align with user preferences across different clinical contexts. To address this challenge, we propose SPA, a segmentation framework that adapts to user preferences. These preferences are modeled using a Gaussian Mixture distribution $p_{\theta}(z)=\sum^{M}_{m=1}\pi_{m}\mathcal{N}(z|\mu_{m},\sigma_{m}^{2})$, where parameters $\theta=\{(\mu_{m},\sigma^{2}_{m},\pi_{m})\}_{m=1}^{M}$ define the preference distribution $p_{\theta}(z)$. Samples are drawn from this distribution as a conditioning signal to capture variations under user preferences  [^1] Preliminary experiments with MRNet [16] and findings in Appendix suggest that while segmentation performance varies across clinicians, each clinician demonstrates consistent results. These observations indicate that modeling interaction behavior can effectively learn user preferences, thereby improving model efficiency.. During training, the preference distribution $p_{\theta}(z)$ captures preference-related segmentation variations. At inference, the model efficiently adapts to user preferences by estimating the maximum likelihood based on user’s (clinician) feedback.
@@
Our goal is to train a general function $f(\cdot,\cdot)$ that can adapt to different user preferences $p$’s through a multi-choice interaction process: $\mathbf{y}=f(\mathbf{x},\mathbf{\Delta_{r_{p}}})$, where $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$ is the input image, and $\mathbf{\Delta_{r_{p}}}$ reflects user interaction, in the form of a selection among multiple ways for correcting the current segmentation. Specifically, the inference of our SPA framework consists of two main steps (Fig. 2): Preference-aware Segmentation (Section 3.2) and Preference Adaption with Human Feedback (Section 3.3). The Preference-aware Segmentation generates multiple valid segmentations to represent image uncertainty and the Preference Adaption with Human Feedback refines these segmentations iteratively to align with specific user preferences. As shown in the upper part of Fig. 2, the model function $g(\cdot,\cdot,\cdot)$ processes three components in the Preference-aware Segmentation: the input image $\mathbf{x}$, interaction embedding $e^{(j)}_{p}$, and $\{\mathbf{z_{n}}^{(j)}\}_{n=1}^{N}$ sampled from the preference distribution $p_{\theta}^{(j)}(z)$ at timestamp $j$. Then, model generates $N$ predictions $\{\mathbf{\hat{y}_{n}}^{(j)}\}_{n=1}^{N}$ to represent uncertainty across individual preferences, where $\mathbf{z_{n}}^{(j)}$’s control the subtle shape based on the user feedback. These predictions are combined into a aggregated preference-aware prediction $\mathbf{\hat{y}_{app}}^{(j)}$ and clustered into $K$ correction proposals $\{\mathbf{\Delta_{r_{k}}}^{(j)}\}_{k=1}^{K}$. The aggregated preference-aware prediction $\mathbf{\hat{y}_{app}}^{(j)}$ is generated through majority vote and then binarized. If the clinician approves with $\mathbf{\hat{y}_{app}}^{(j)}$, it will be used as the final segmentation and the iteration terminates. Otherwise, the framework uses Preference Adaption with Human Feedback step for further adaption, as shown in the lower part of Fig. 2: The clinician selects the proposal $\mathbf{\Delta_{r_{p}}}^{(j)}$ that best aligns with their preference $p$. The preference distribution $p_{\theta}^{(j)}(z)$ is then updated to $p_{\theta}^{(j+1)}(z)$ based on this feedback. This iterative process continues until the clinician satisfies with the prediction.
By doing so, SPA gain three advantages over previous methods: First, it explicitly models diverse human preferences incorporated in multi-annotator datasets with a preference distribution $p_{\theta}(z)$ in training time, increasing the chance for a preferred segmentation to be reachable at inference time. Second, it improves efficiency by adapting predictions with fewer interaction rounds than existing interactive segmentation methods. Third, it reduces clinician’s effort at each user interaction by replacing pixel-level pinpoint with a simpler multi-choice selection approach.
@@ 3.2 Preference-aware Segmentation
As shown in the upper part of Fig. 2, our Preference-aware Segmentation step involves the synergy of an image encoder and a mask decoder. Given an input image $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$, we first obtain a general image embedding $\mathbf{e_{x}}^{(j)}\in\mathbb{R}^{L\times{\frac{H}{P}}\times{\frac{W}{P}}}$ at time $j$, where $L$ represents the output channels and $P$ is the Vision Transformer (ViT) patch size. This embedding is generated using a pre-trained ViT, following the architecture described in MAE [12]. Additionally, we sample $N$ preference conditions $\{\mathbf{z_{n}}^{(j)}\}_{n=1}^{N}$ from the preference distribution $p_{\theta}^{(j)}(z)$. Each $\mathbf{z_{n}}^{(j)}\in\mathbb{R}^{L\times{\frac{H}{P}}\times{\frac{W}{P}}}$ is concatenated with the general image embedding $\mathbf{e_{x}}^{(j)}$, and then processed through convolutional layers and ReLU activations to produce a set of preference-aware image embeddings $\{\mathbf{e_{x_{n}}}^{(j)}\}_{n=1}^{N}$, where each $\mathbf{e_{x_{n}}}^{(j)}\in\mathbb{R}^{L\times{\frac{H}{P}}\times{\frac{W}{P}}}$.
@@
$\mathbf{e_{x_{n}}}^{(j)}=ReLU(Conv(\mathbf{e_{x}}^{(j)}\oplus\mathbf{z_{n}}^{(j)})),n=1,2,...,N$ (1)
@@
To predict dense semantic masks from the preference-aware image embeddings $\{\mathbf{e_{x_{n}}}^{(j)}\}_{n=1}^{N}$ and the interaction embedding $e_{p}^{(j)}$ (Section 3.3), we employ a mask decoder that follows the same architecture as SAM [18]. This decoder predicts $N$ segmentation masks $\{\mathbf{\hat{y}_{n}}^{(j)}\}_{n=1}^{N}$ for the input image $\mathbf{x}$ at timestamp $j$. The aggregated preference-aware prediction $\mathbf{\hat{y}_{app}}^{(j)}$ is generated by applying a majority vote on the predictions $\{\mathbf{\hat{y}_{n}}^{(j)}\}_{n=1}^{N}$, followed by binarization.

### v2

@@ Abstract
Medical image segmentation data inherently contain uncertainty. This can stem from both imperfect image quality and variability in labeling preferences on ambiguous pixels, which depend on annotator expertise and the clinical context of the annotations. For instance, a boundary pixel might be labeled as tumor in diagnosis to avoid under-estimation of severity, but as normal tissue in radiotherapy to prevent damage to sensitive structures. As segmentation preferences vary across downstream applications, it is often desirable for an image segmentation model to offer user-adaptable predictions rather than a fixed output. While prior uncertainty-aware and interactive methods offer adaptability, they are inefficient at test time: uncertainty-aware models require users to choose from numerous similar outputs, while interactive models demand significant user input through click or box prompts to refine segmentation. To address these challenges, we propose SPA, a new Segmentation Preference Alignment framework that efficiently adapts to diverse test-time preferences with minimal human interaction. By presenting users with a select few, distinct segmentation candidates that best capture uncertainties, it reduces the user workload to reach the preferred segmentation. To accommodate user preference, we introduce a probabilistic mechanism that leverages user feedback to adapt a model’s segmentation preference. The proposed framework is evaluated on several medical image segmentation tasks: color fundus images, lung lesion and kidney CT scans, MRI scans of brain and prostate. SPA shows 1) a significant reduction in user time and effort compared to existing interactive segmentation approaches, 2) strong adaptability based on human feedback, and 3) state-of-the-art image segmentation performance across different imaging modalities and semantic labels. Our code is publicly available here: https://github.com/SuperMedIntel/SPA.
@@ 1 Introduction
Deep learning-based medical image segmentation has achieved remarkable progress over the past decade [27, 31, 5]. However, existing approaches often fail when applied to real-world clinical scenarios. A critical challenge is how to handle the inherent uncertainties in medical images [13]. A single medical image may have multiple different valid segmentation results, depending on the labeling criteria for a specific medical context. For example, in glioma detection from brain CT scans, it is often preferred to include surrounding tissue rather than risk missing part of the tumor [2]. In contrast, in radiation therapy for low-grade glioma, undersegmentation is preferred to protect sensitive brain tissue from excessive radiation damage [26]. Therefore, it is essential to develop adaptive methodologies that align segmentation uncertainties with specific labeling preferences for different clinical needs.
@@
Figure 1: (a) Existing uncertainty-aware models require users to choose from numerous similar-looking candidates, making the process labor-intensive and time-consuming. (b) Interactive segmentation models lack the ability to incorporate image uncertainty and rely heavily on pixel-level user inputs, which require a substantial amount of time and effort. (c) Our uncertainty-aware interactive segmentation model, SPA, efficiently achieves segmentations whose decisions on uncertain pixels are aligned with users preferences. This is achieved by modeling uncertainties and human interactions. At inference time, users are presented with one recommended prediction and a few representative segmentations that capture uncertainty, allowing users to select the one best aligned with their clinical needs. If the user is unsatisfied with the recommended prediction, the model learns from the users’ selections, adapts itself, and presents users a new set of representative segmentations. Our approach minimizes user interactions and eliminates the need for painstaking pixel-wise adjustments compared to conventional interactive segmentation models.
@@
Existing uncertainty-aware approaches [17, 28, 3, 22] represent segmentation uncertainty to users by generating numerous stochastic predictions (Fig. 1 (a)). However, users often have to carefully choose from countless similar-looking candidates, making the process time-consuming. Furthermore, since these models cannot incorporate human feedback for adjustments, there is a risk that none of the predictions be satisfying.
@@
Incorporating human feedback into segmentation, often in the form of visual prompts (such as clicks and bounding boxes) has shown promise [30, 6, 24, 18]. However, most such interactive segmentation approaches do not incorporate uncertainty. Additionally, existing visual prompts are often cumbersome to attain in the real world, as they require pixel-level interaction input from a user (Fig. 1 (b)).
@@
To address the above challenges, we propose to present users a small number of distinct segmentations to represent uncertainty. It allows users to select their preferred option, simplifying interactive refinement into a straightforward multiple-choice selection. In this paper, we introduce SPA, a new approach for efficient Segmentation Preference Alignment with uncertainty in medical image segmentation (Fig. 1). SPA presents image uncertainties by generating multiple segmentations. Instead of providing users with numerous similar-looking predictions as in conventional uncertainty-aware segmentations, our model offers one recommended prediction and four representative segmentation candidates per iteration. Once the user selects an option, SPA rapidly adapts through iterative refinement, aligning efficiently with the user’s segmentation preference in only a few iterations. Our experiment demonstrates that a user can segment 35% more images with 39% fewer iterations compared to previous interactive models. This highlights SPA’s potential for real-world clinical applications. In summary, our contributions are:
  • We introduce SPA, a new segmentation framework that can align to users’ decision preferences toward ambiguous/uncertain pixels at inference time. Such preferences often vary with clinical contexts, which existing segmentation models cannot adapt to efficiently.
  • We propose to model uncertainties in pixel-wise predictions under diverse preferences as a parameterized latent distribution. This modeling enables rapid adaptation to user preferences with minimal iterations at inference time.
  • We develop a multi-choice interaction mechanism to receive user preferences at inference time, providing a more friendly and effortless user experience than existing works.
@@
  • We compare SPA with deterministic, uncertainty-aware, and interactive SOTA methods. SPA achieves superior results measured by Dice Score for three multi-clinician annotated datasets (REFUGE2, LIDC-IDRI and QUBIQ). SPA also consistently outperforms interactive methods in terms of requiring fewer rounds of user interactions.
@@
Figure 2: Overall framework of SPA. The inference process comprises of two steps: Preference-aware Segmentation and Preference Adaptation with Human Feedback. At iteration $j$, SPA takes the input image $\mathbf{x}$, an interaction embedding $e_{u}^{(j)}$, and latent variables $\{\mathbf{z_{n}}^{(j)}\}_{n=1}^{N}$ drawn from the preference distribution $p^{(j)}_{\theta}(z)$ to generate $N$ segmentation predictions. These predictions are then combined into an aggregated preference-aware prediction $\mathbf{{y}_{app}}^{(j)}$. If the user is not satisfied with $\mathbf{{y}_{app}}^{(j)}$, SPA generates $K$ representative segmentation candidates $\{\mathbf{r_{k}}^{(j)}\}_{k=1}^{K}$ in the adaptation step. The user selects the preferred segmentation candidate $\mathbf{r_{u}}^{(j)}$. The preference distribution is then updated to $p^{(j+1)}_{\theta}(z)$ based on this choice. This process iterates until the segmentation meets clinical satisfaction.
@@ 2 Related Work
Uncertainty-aware Medical Segmentation Uncertainty in medical images cannot be reduced by adding more data or using more complex models [19]. Techniques such as model ensembling, label sampling [15], and multi-head strategies [11] attempt to address this uncertainty by generating a range of potential predictions that reflect different user preferences [36, 14]. Probabilistic segmentation methods, including ProbUNet [20], PhiSeg [3], CM-Net [34] and MRNet [16] explicitly model the posterior distribution of parameters or predictions to capture uncertainty. However, these approaches often generate multiple predictions, requiring users to review each individually. Moreover, the predictions may not perfectly match the specific clinical context, and some techniques rely on prior knowledge of the expertise of a clinician, complicating adoption into clinical practice. In contrast, a more efficient approach is to incorporate human interactions, allowing users to refine segmentations through direct interaction to better align with the clinical context.
Interactive Medical Segmentation Interactive segmentation is an iterative process where automated segmentation results are refined through user input until they reach a desired output. Previous methods [32, 29, 18] have modified predictions based on user interaction at the pixel level, with some achieving success in the medical domain [9, 25, 33, 35]. However, these interactive models often do not address the inherent uncertainty in medical images. Additionally, they may require numerous iterations to align with the user preference for a specific clinical context, given the existence of multiple valid segmentations. An interactive model capable of incorporating image uncertainties and learning from user interactions would reduce the number of iterations required to achieve a desired result.
@@ 3 Methodology
3.1 Model user preferences as latent distributions
We assume that each user’s decisions on uncertain/ambiguous pixels, based on their preference and stemming from different medical contexts, can be modeled as i.i.d. samples drawn from a parameterized distribution. To represent this user preferences for different clinical contexts, we model this preference distribution $p_{\theta}(z)$ as a mixture of $M$ Gaussian distributions:
where $\mathcal{N}(z\mid\mu_{m},\sigma_{m}^{2})$ denotes the Gaussian component for user $m$, with mean $\mu_{m}$ and variance $\sigma_{m}^{2}$. The mixture weights $\pi_{m}$ satisfy $\pi_{m}>0$ and $\sum_{m=1}^{M}\pi_{m}=1$. Each component in this mixture model reflects a specific user segmentation preference, enabling the framework to generate segmentation predictions that capture discrepancies across users. In our segmentation framework, $z\sim p_{\theta}(z)$ is implemented as a conditioning signal injected into a neural network, predicating the subtle preference-dependent variability in segmentation outputs.
Modeling user preferences as parameterized distributions is highly adaptable, as the distribution $p_{\theta}(z)$ can be iteratively updated via maximum likelihood based on new observations of user interactions at inference time  [^1] We provide a theoretical proof in the appendix to demonstrate that, given sufficient interactions from a specific user $u$, the preference distribution $p_{\theta}(z)$ will converge to the user’s personalized distribution $\mathcal{N}(\mu_{u},\sigma_{u}^{2})$..
@@ 3.1 Model user preferences as latent distributions
We integrate the preference distribution $p_{\theta}(z)$ into the segmentation workflow, as follows. In each iteration, we sample latent variables $\{\mathbf{z_{n}}\}_{n=1}^{N}$ from $p_{\theta}(z)$ to capture variations under user preferences  [^2] Preliminary experiments with MRNet [16] given in the appendix suggest that while segmentation performance varies between users, each user demonstrates consistent results. These observations indicate that modeling interaction behavior can effectively learn user preferences, thereby improving modeling efficiency.. During training, the model learns $p_{\theta}(z)$ through user feedback simulated from annotated segmentations from multiple users. At inference, the model can efficiently adapt to user preferences by estimating the maximum likelihood based on the feedback of a user.
3.2 Overall Workflow
We propose SPA, which learns to model human preferences by preference distribution $p_{\theta}(z)$ during training, enabling it to efficiently model new user preferences at inference through interaction. Our framework also simplifies and improves the robustness of interaction by replacing pixel-level clicking with multi-choice selection.
Formally, our goal is to learn a general function $f(\cdot,\cdot)$ that can adapt to different user preferences $u$’s through an interaction process: $\mathbf{y}=f(\mathbf{x},\mathbf{r_{u}})$, where $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$ is the input image, and $\mathbf{r_{u}}$ reflects human interaction. The training details are provided in Section 3.5 and Algorithm 1.
Specifically, our SPA framework consists of two main modules (Fig. 2): Preference-aware Segmentation (Section 3.3) and Preference Adaption with Human Feedback (Section 3.4). Our Preference-aware Segmentation generates multiple valid segmentations to represent image uncertainty (oftentimes aleatoric) and the Preference Adaption with Human Feedback aligns these segmentations iteratively with specific user preferences.
@@ 3.2 Overall Workflow
In inference, as shown in the upper part of Fig. 2, given a raw image $\mathbf{x}$, Preference-aware Segmentation generates $N$ segmentation predictions $\{\mathbf{{y}_{n}}^{(j)}\}_{n=1}^{N}$, conditioned by the interaction embedding $e^{(j)}_{u}$ representing the user selection, and latent variables $\{\mathbf{z_{n}}^{(j)}\}_{n=1}^{N}$ sampled from the preference distribution $p_{\theta}^{(j)}(z)$ at iteration $j$. We use the user’s first point prompt to produce $e^{0}_{u}$ and $p_{\theta}^{0}(z)$ for the initialization. Predictions $\{\mathbf{{y}_{n}}^{(j)}\}_{n=1}^{N}$ reflect uncertainty from individual subtle preferences, which is controlled by $\mathbf{z_{n}}^{(j)}$. These predictions are combined into one aggregated preference-aware prediction $\mathbf{{y}_{app}}^{(j)}$ and $K$ Representative Segmentation Candidates $\{\mathbf{r_{k}}^{(j)}\}_{k=1}^{K}$. If the user approves $\mathbf{{y}_{app}}^{(j)}$, it will be used as the final segmentation and the iteration ends. Otherwise, the framework uses the Preference Adaption with Human Feedback step for further refinement, as shown in the lower part of Fig. 2: The user then selects the segmentation $\mathbf{r_{u}}^{(j)}$ that best aligns with their preference. The preference distribution $p_{\theta}^{(j)}(z)$ is further updated to $p_{\theta}^{(j+1)}(z)$ based on this feedback. This iterative process continues until the user is satisfied with the segmentation.
@@ 3.3 Preference-aware Segmentation
As shown in the upper part of Fig. 2, our Preference-aware Segmentation step involves an image encoder and a mask decoder. Given an input image $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$, we first obtain a general image embedding $\mathbf{e_{x}}^{(j)}\in\mathbb{R}^{L\times{\frac{H}{P}}\times{\frac{W}{P}}}$ for iteration $j$, where $L$ represents the output channels and $P$ is the Vision Transformer (ViT) patch size. This embedding is generated using a pre-trained ViT [12]. Additionally, we sample $N$ preference conditions $\{\mathbf{z_{n}}^{(j)}\}_{n=1}^{N}$ from the preference distribution $p_{\theta}^{(j)}(z)$. Each $\mathbf{z_{n}}^{(j)}\in\mathbb{R}^{L\times{\frac{H}{P}}\times{\frac{W}{P}}}$ is concatenated with the general image embedding $\mathbf{e_{x}}^{(j)}$, and then processed through three convolutional layers and ReLU activations to produce a set of preference-aware image embeddings $\{\mathbf{e_{x_{n}}}^{(j)}\}_{n=1}^{N}$, where each $\mathbf{e_{x_{n}}}^{(j)}\in\mathbb{R}^{L\times{\frac{H}{P}}\times{\frac{W}{P}}}$.
Input: Image $\mathbf{x}$, user-preferred representative segmentation candidate $\mathbf{r_{u}}$
Output: Preference-aware Segmentation $\mathbf{y_{app}}$
/* Definitions */
Preference Distribution Generator: $E_{\theta_{\mu_{m};\sigma_{m};\pi_{m}}}(\cdot)$ generating $\{(\mu_{m},\sigma_{m}^{2},\pi_{m})\}_{m=1}^{M}$ to construct $p_{\theta}$; Interaction Encoder: $E_{I}(\cdot)$; Number of samples: $N$; Number of representative segmentation candidates: $K$;
Function GenMask($\mathbf{e_{x}}$, $p_{\theta}$, ${e_{u}}$):
@@
       Sample $n$ embeddings: $\mathbf{z_{n}}\sim p_{\theta}$, then   $\mathbf{e_{x_{n}}}=\text{ReLU}(\text{Conv}(\mathbf{e_{x}}\oplus\mathbf{z_{n}}))$;
       $\mathbf{y_{n}}=\text{Decoder}(\mathbf{e_{x_{n}}},{e_{u}})$;
       $\mathbf{y_{app}}={\mathds{1}_{0.5}}\left(\frac{1}{N}\sum_{n=1}^{N}\mathbf{y_{n}}\right)$;
       return $\mathbf{y_{app}},\{\mathbf{y_{n}}\}_{n=1}^{N}$;
/* Training Loop */
for t = 1 to MAX_TRAINING_UPDATES do
       Initialize $\mathbf{y}$, $p_{\theta}$, $\mathbf{r_{u}}$ as described in Section 3.2
      for j = 1 to MAX_USER_ITERATIONS do
             Extract: $\mathbf{e_{x}}=\text{ViT}(\mathbf{x})$, ${e_{u}}=E_{I}(\mathbf{r_{u}})$;
            Call $\mathbf{y_{app}^{\text{ori}}},\_=\textnormal{{GenMask}}(\mathbf{e_{x}},p_{\theta},{e_{u}})$;
            /* Update Preference Distribution Generator */
             Update mean and variance: $\theta^{{}^{\prime}}_{\mu_{m};\sigma_{m}}=\theta_{\mu_{m};\sigma_{m}}+\alpha\nabla_{\theta_{\mu_{m}};\sigma_{m}}\mathcal{L}_{\text{CE}}(\mathbf{y^{ori}_{app}},\mathbf{y})$;
            Compute GMM responsibilities (E-step): $\pi_{m}^{gt}=\frac{\pi_{m}p(y\mid\mu_{m},\sigma_{m}^{2})}{\sum_{m=1}^{M}\pi_{m}p(y\mid\mu_{m},\sigma_{m}^{2})}$;
            Update GMM weights: $\theta^{{}^{\prime}}_{\pi_{m}}=\theta_{\pi_{m}}+\alpha\nabla_{\theta_{\pi_{m}}}\mathcal{L}_{\text{MSE}}(\pi_{m}^{gt},\pi_{m})$;
            Call $\mathbf{y_{app}^{\text{gmm}}},\_=\textnormal{{GenMask}}(\mathbf{e_{x}},p_{\theta}(\theta^{{}^{\prime}}_{\mu_{m};\sigma_{m}};\theta^{{}^{\prime}}_{\pi_{m}}),{e_{u}})$;
            /* Segmentation Update */
             Update segmentation and interaction encoders: $\theta_{PSeg}^{{}^{\prime}}=\theta_{PSeg}+\alpha\nabla_{\theta_{PSeg}}\mathcal{L}_{\text{CE}}(\mathbf{y_{app}^{\text{gmm}}},\mathbf{y})$, $\theta_{E_{I}}^{{}^{\prime}}=\theta_{E_{I}}+\alpha\nabla_{\theta_{E_{I}}}\mathcal{L}_{\text{CE}}(\mathbf{y_{app}^{\text{gmm}}},\mathbf{y})$;
            Call $\mathbf{y_{app}^{\text{fin}}},\mathbf{\{y_{n}^{\text{fin}}\}}_{n=1}^{N}=\textnormal{{GenMask}}(\mathbf{e_{x}^{\prime}},p_{\theta}(\theta^{{}^{\prime}}_{\mu_{m};\sigma_{m}};\theta^{{}^{\prime}}_{\pi_{m}}),{e_{u}^{\prime}})$;
            /* Simulate User Selection for the Next Iteration */
             Get $K$ Representative Candidates: $\{\mathbf{{r_{k}}}\}_{k=1}^{K}=\text{KMeans}(\{\mathbf{y_{n}^{\text{fin}}}\}_{n=1}^{N})$;
            Select preferred representative candidate: $\mathbf{r_{u}}=\arg\min_{\mathbf{r_{k}}}\|\mathbf{r_{k}}-\mathbf{y_{app}^{\text{fin}}}\|_{2}$;
Algorithm 1 SPA Training Process
@@
We then employ a mask decoder to predict segmentation masks $\{\mathbf{{y}_{n}}^{(j)}\}_{n=1}^{N}$ from $\{\mathbf{e_{x_{n}}}^{(j)}\}_{n=1}^{N}$ and the interaction embedding $e_{u}^{(j)}$ (Section 3.4). The decoder follows the same architecture as SAM [18]. The final preference-aware prediction is aggregated as $\mathbf{y_{app}}^{(j)}={\mathds{1}_{0.5}}\left(\frac{1}{N}\sum_{n=1}^{N}\mathbf{y_{n}}^{(j)}\right)$.